Gas systems project

Europe/Zurich
Beatrice Mandelli (CERN)
Videoconference
Gas systems project
Zoom Meeting ID
67695456084
Host
Beatrice Mandelli
Useful links
Join via phone
Zoom URL

GAS M&O meeting 17 September 2021

Present: Beatrice, Roberto, Kacper, Michal, Fred, Patrick, Andrea, Mattia, Fabio, Loris

Indico: https://indico.cern.ch/event/1076617/

 

ALICE #no elog

TPC

To remember

  • new MFC to install
  • pump regulation valve noisy since the restart (~10%) —> to check old PID

From previous meeting (05/10)

  • Krohne: only present one readying for the full system (distribution is just one channel) —> Leonardo gave to Michal the formula (15/10) —> Michal to implement

TRD

  • purifier Xe recuperation humidity at -10 C —> Gianluca to check if is ok

To remember:

  • distribution module: one big difference between PT and pressure gauge: to check which one is wrong
  • T sensor in pump: Patrick said he can do but it takes time
  • difficult to restart: need to work on PID of pump and distribution when correct gas present
  • problem of noise in the read-out of Krohne to investigate
  • Brown bottle to fill: tested in another bottle and from the test we know now that it has to be done all manual --> LPh to check with Chilo when to do
  • disk de rupture: they work —> LPh has to contact the safety

HMP

From previous meeting (03/08):

  • Distribution rack in alarm —> LPh to check

MTR

  • MTR running in closed loop with correct gas mixture, difficult to restart the system:
    • humidifier MFCs didn’t open
    • mixer MFC very very slow to change setpoint
    • to remember to change back-up gas from Ar to Freon in distribution rack (manual valve), P&ID not update in distribution rack
    • very low flow for VAISALA humidifier: it didn’t react well —> flow increased
  • at start of the mixer, SF6 concentration is not correct —> Michal/Gianluca to investigate
  • back-up to check
  • pressure sensor on the chamber doesn’t work —> LPh
  • Tescom of the Zimmerli of the exhaust broken —> LPh to change
  • Zimmerli of exhaust and mixer don’t have the remote control —> LPh to implement
  • MFCs of humidifier and exhaust in error (but it seems they work) —> Michal/Patrick to investigate

From previous meeting (22/01)

  • pump regulation to check as it doesn’t work well at the start of the system
  • to connect dummy chamber electrically —> LPh

To remember

  • leak test and maintenance in the distribution —> LPh/Mattia
  • No dummy chamber present in software —> Michal

CPV

  • found in Run but with alarms disabled

TOF

To remember

MCH

To remember

  • still some work in PID tuning

PGS

To remember

  • SF6 PGS is scheduled in 2022 —> LPh to check with David if it can be done this year

 

ATLAS #no elog

MDT

To remember:

  • found 60 ppm of offset in Vaisala (but within the specification): new calibration to do (bottle ordered)
  • list of pressure sensors replaced in the last years —> Kacper
  • some spikes in flowcells after the upgrade, spikes also up to 3000 l/h so it is not a real flow spike —> Gianluca to prepare few examples
  • Rack 73: change of ELMB and now ok. Michal to check the ELMB in the lab
  • problem in one VAISALA reading but it is not the sensor —> to investigate

RPC

  • UPS in SGX1 ready —> to schedule installation —> Beatrice/Kacper to check with Patrick
  • rack 61, 62, 63, 67 ON
  • Beatrice to check for UPS in underground

To remember:

  • purifier 3: missing MFC for conditioning
  • UPS for ATLAS RPC system
  • changed motherboard/ELMB in rack 63 because of channel 2 in error (it was already changed in 2020)—> Michal to check motherboard/ELMB removed

TGC

  • gas system on
  • rack 63 flowcells removed for calibration (like rack 67 where there was a problem in calibration)
  • exhaust modification on-going
  • studies on-going for flowcell/elmb/motherboards

TRT

From previous meeting (23/10)

  • Krohne new formula prepared by Leonardo —> Michal to implement
  • Krone exhaust not working properly: to check if it is the krohne or our electrical part. Patrick to try to check the electrical part
  • Kacper to check for regeneration of purifier material

MM

  • system on, detectors connected
  • rack 61: to check electrically to connect to MM gas system —> Patrick/Kacper
  • to check for DSS signal —> Beatrice/Kacper/Patrick

 

CMS #3532

RPC

  • flowcells of rack 77 have a deviation of 75%. Michal recalibrated at 60 l/h —> to check if this value is ok (maximum flow foreseen by RPC team: 30 l/h)
  • Andrea removed lateral doors for UX racks to have access to flexible pipes with B-field. Andrea will put protection for bubblers
  • high flow test next week

To remember:

  • low flow MFC (1200 l/h) received —> to plan installation
  • iC4H10 problem: still 0.1% difference—> Mara to check with the company:
    • according to company calibration should be performed once/month. coherent with discrepancy we see, develops after ~ 4-6 weeks. Calibration has to be done manually, 5% isobutane bottle is flammable, cannot stay permanently connected in rack (checked with R. Perruzza)
  • Group regulation valves that are at 85% for group 1 and 5 —> Andrea/Mara to check

CSC

  • zimmerli in parallel to MFC of exhaust didn’t open when recovery stop —> to do some tests
  • pump membrane changed
  • canbus cable to be replaced between rack 65 and service cavern —> Mara/Andrea

To remember to do

  • Difference in PT5005-PT5010: is it in the purifier? --> Mara to check
  • To put scale sticker on all bubblers —> Andrea-
  • new Von Rohr regulation valve to test

DT

  • changed PLC power supply and changed H2O analysers, values look ok
  • zimmerli of exhaust not mounted because it still doesn’t work —> Andrea to bring to Meyrin to do some test with Mara
  • maintenance mixer done without leak test: no enough time
  • for next time: to ask several days of stop for doing the maintenance

To remember

  • test on the exhaust module since when there is a start after a stop we loose ~0.1-0.2 bar: it looks like regulation valves in distribution open too fast. PID changed, to do other tests

GEM

  • Changed PLC power supply
  • safety valve at 75 mbar installed in GE-1

To remember:

  • safety valves: 50 mbar in + side and 70 mbar in - side. To change in positive side

TRACKER

  • MFM to install, waiting green light form users

 

 

LHCb #3539

RICH2

  • open to Air with Air compressor on connected at one of the input
  • alarm on pressure chamber activated
  • CO2 back-up valve closed

To remember

  • IR analysis: to add pump for recycling the gas —> Patrick/Mara
  • to investigate stop of RICH2: maybe related to profibus? module WAGO changed
  • Leak test of purifier on-going: done but not report
  • New leak appeared: probably at detector level. Patrick is checking with Christoph
    • Detector people changed some connectors of cables at the detector level
  • PT5005 changed but still spikes visible: to think about
  • problem of high injection of CO2 from the mixer
    • probably due to purifier (leak in column A + not saturated in CO2)
    • leak already present in 2018 (elog #2886)

MWPC

  • start CF4 on Monday
  • motherboard changed in racks 63 and 66

From previous meeting

  • SG-UX pipe to test
  • check of flowcells reading and check with Oleg if he does some leak test —> Mara/Patrick
    • impossible to do leak test at detector level unless we know exactly what distribution section to test no leak search was ever performed on their side as it is very difficult  side A can be accessed in autumn but they cannot test all chambers
  • test sending the gas at the exhaust: mixer 77 l/h, exhaust 66 l/h

Reminder

  • found leak, probably on distribution: new test and no leak found

ANALYSIS CHAIN 3

Reminder

  • leaks in electro-valves —> to change all

SciFi

  • it was on with alarms disabled and with wrong levels
  • next week several test to check status and alarm level
  • the 2 vaisala have a discrepancy of 8 C —> Mara to check

PGS

  • Mara contacted David: all ok to start CF4

 

NON-LHC

  • All ok
  • Some valves in force mode in LINAC —> Fabio/Roberto to put comment
  • NA62 RICH analysis was off (without gas) since the PLC crash

Reminder:

  • LINAC4: to buy a spare PLC Siemens —> Michal next year
  • NA62: to check if we need to buy only the new PLC or also the I/O cards —> Michal

 

SOFTWARE

  • new ROG will become compulsory in the future. Michal prepared few slides: end of the year we’ll not have anymore our std system —> everybody to try asap
  • PLC planning until March 2022 shown

From previous meeting

  • new monitoring page “Exhaust++” done for ATLAS MDT explained by Gianluca. Everybody to look and give comments
    • next step: to do a similar page for one gas system for each experiment
  • To check for having warning in difference between PT5011 and PT5005
  • New feature for alarm when system is OFF: to provide to Michal the list of alarms we want to have when system are OFF —> Mara/Gianluca with experiment technical responsible by week 5
  • Profibus/canbus monitoring: Michal shows where we have and where it could be useful to have (based on history of problems in the past). To discuss if regeneration of code is worst for these system
  • We need to define how to proceed for the classification of warnings and alarms
  • Slide on the modifications asked by Schniedr with pictures and time needed to do modification
  • To fix the StepWS problem asap otherwise difficult to have history in case of problems —> Michal/Gianluca
  • Patrick will prepare a box with all spare material for PLC. Box will be in electrical workshop
  • For system with old PLCs: Michal to add in the rack an old spare PLC
  • For TFC: Michal to ask to Majec for a spare IO card (special one for TFC)

 

PIQUET

  • Discussion of alarms week 36-37:
    • ALICE MTR: Alarm: purifier module 1 not in Nominal Run. 
    • ALICE PGS: HH flow for N2, not understood —> LPh to check when back
    • ALICE PGS: HH flow for CO2 due to the use of CO2 by TPC purifier —> to change flow for next time —> LPh/Gianluca
    • CMS CSC: alarm on regulation valve of pump due to high pressure in the loop because CF4 recovery stopped but zimmerli of exhaust didn’t open. CF4 recovery without power (external cause)
    • During routine check:
      • Found LHCb SciFi and ALICE CPV in Run with alarms disabled
      • Some alarm forgot blocked in some system
      • ALICE Analysis Chain 3 in force mode with valve of TOF mixer open to air (lost of ~10-20 l/h of gas) since 28/05

 

GENERAL

  • Beatrice show dates for closing of caverns in 2022

From previous meeting

  • Found problem of PROFIBUS MDT: termination of PROFIBUS PA on the connector was wrong —> to change in ALL gas systems
  • Aging test priority
    • To do: Airliquide pressure reducer, IR iC4H10 BINOS (they do aging —> LPh to check for modification)
There are minutes attached to this event. Show them.
    • 2:00 PM 2:05 PM
      Piquet interventions 5m

      http://eloggas.cern.ch/GAS_PIQUET/

      Speakers: Andrea D'Auria (CERN) , Beatrice Mandelli (CERN) , Frederic Merlet (CERN) , Louis-Philippe De Menezes (CERN) , Patrick Carrie (CERN) , Kacper Jerzy Kapusniak (CERN) , Michal Zbigniew Zimny (CERN) , Gianluca Rigoletti (Universite Claude Bernard Lyon I (FR))
    • 2:05 PM 2:15 PM
      M&O and project organization 10m
      Speakers: Beatrice Mandelli (CERN) , Roberto Guida (CERN)

      CMS closed: Sept 22nd

      • Cavern closed: Sept 24th

    • 2:15 PM 2:25 PM
      New ROG system and PLC planning 10m
      Speaker: Michal Zbigniew Zimny (CERN)

      LS2 upgrades planning:

      https://readthedocs.web.cern.ch/display/GCSD/Provisional+upgrades+planning

      CLOUD - firs version ready with IOs

      ND280 - 11.10.2021

      CMS RPC - 09.11.2021

      CMS AUX - 23.11.2021 - still to confirm if possible to do on 15.11.2021

      LHCb RI1 - 3.12.2021

      LHCb Flushing - end of 2021

      NA62 - 25.01.2022

      CMS CSC - 4.02.2022

      ALICE AUX - postpone to one Technical Stop during Run 3

      LHCb MWPC - if it is possible I would like to do in February 2022, otherwise we will schedule for one TS of Run3

    • 2:25 PM 2:30 PM
      Activities on-going 5m
      Speakers: Beatrice Mandelli (CERN) , Frederic Merlet (CERN)
    • 2:30 PM 2:50 PM
      Plan of LS2 activities 20m
      Speakers: Beatrice Mandelli (CERN) , Frederic Merlet (CERN) , Roberto Guida (CERN)
    • 2:50 PM 3:10 PM
      NA61 and ND280 update 20m
      Speakers: Andrea D'Auria (CERN) , Fabio Bordon (CERN) , Frederic Merlet (CERN) , Kacper Jerzy Kapusniak (CERN) , Mattia Busato (Institut National Polytechnique de Grenoble (FR)) , Patrick Carrie (CERN)
    • 3:10 PM 3:20 PM
      CLOUD update 10m
      Speakers: Louis-Philippe De Menezes (CERN) , Patrick Carrie (CERN)
    • 3:20 PM 4:45 PM
      Gas systems for the LHC experiments
      • 3:35 PM
        ALICE 5m
        Speakers: Gianluca Rigoletti (Universite Claude Bernard Lyon I (FR)) , Louis-Philippe De Menezes (CERN)
      • 3:40 PM
        ATLAS 5m
        Speakers: Gianluca Rigoletti (Universite Claude Bernard Lyon I (FR)) , Kacper Jerzy Kapusniak (CERN)
      • 3:45 PM
        CMS 5m
        Speakers: Andrea D'Auria (CERN) , Mara Corbetta (AGH University of Science and Technology (PL))
      • 3:50 PM
        LHCb 5m
        Speakers: Mara Corbetta (AGH University of Science and Technology (PL)) , Patrick Carrie (CERN)
      • 3:55 PM
        Not-LHC experiments 5m
        Speaker: Fabio Bordon (CERN)
      • 4:00 PM
        Software 10m
        Speakers: Gianluca Rigoletti (Universite Claude Bernard Lyon I (FR)) , Michal Zbigniew Zimny (CERN)
      • 4:10 PM
        Gas web-page 5m

        http://test-gasmonitoring.web.cern.ch
        https://ep-dt-gasanalysis.web.cern.ch

        Speaker: Beatrice Mandelli (CERN)
      • 4:15 PM
        Gas analysis: standard H2O/O2 5m
        Speaker: Louis-Philippe De Menezes (CERN)
    • 4:45 PM 5:35 PM
      Gas System developments
      • 4:45 PM
        Gas System Test-Bench 10m
        Speaker: Fabio Bordon (CERN)
      • 4:55 PM
        CMS CF4 recovery 10m
        Speakers: Fabio Bordon (CERN) , Mara Corbetta (AGH University of Science and Technology (PL))
      • 5:05 PM
        LHCb RICH2 CF4 recuperation 10m
        Speaker: Mara Corbetta (AGH University of Science and Technology (PL))
      • 5:15 PM
        R134a Recuperation 10m
        Speakers: Gianluca Rigoletti (Universite Claude Bernard Lyon I (FR)) , Mara Corbetta (AGH University of Science and Technology (PL))
      • 5:25 PM
        AoB 10m
        Speakers: Fabio Bordon (CERN) , Gianluca Rigoletti (Universite Claude Bernard Lyon I (FR)) , Mara Corbetta (AGH University of Science and Technology (PL))
    • 5:35 PM 5:45 PM
      FSU & Jobs status 10m
      Speaker: Frederic Merlet (CERN)
    • 5:45 PM 5:55 PM
      AoB 10m
      Speakers: Andrea D'Auria (CERN) , Beatrice Mandelli (CERN) , Frederic Merlet (CERN) , Kacper Jerzy Kapusniak (CERN) , Louis-Philippe De Menezes (CERN) , Patrick Carrie (CERN) , Roberto Guida (CERN)